OpenStudy (anonymous):

add/simplify 5x*x^2-49 +x/x-7

OpenStudy (anonymous):

5x^3-55 x/x=1 -49-7=-56 -56+1=55 x*x^2=x^3

OpenStudy (anonymous):

this is unless you miswrote the equation, which it looks like you did

jhonyy9 (jhonyy9):

please us parantheses